OpenStudy (rianiscool):

need help asap will give a medal Enter the answer. 6. Sharon is the youngest in her family. Her sister Anne is 3 years older than she is. Her brother Tom is 5 years older than Anne. The sum of their ages is 62. How old is Sharon?

OpenStudy (anonymous):

this is a hard one

OpenStudy (rianiscool):

?

OpenStudy (akashdeepdeb):

Let Sharon's age = x Anne = x-3 [Because, Anne is younger] Tom = (x-3) + 5 = x+2 [5 years older than Anne] Sum = x + x-3 + x+2 = 3x - 1 = 62 3x-1 = 62 [Solve this, and you'll get Sharon's age.]